# CSPCL: Category Semantic Prior Contrastive Learning for Deformable DETR-Based Prohibited Item Detectors

### v1

@@ Abstract
Prohibited item detection based on X-ray images is one of the most effective security inspection methods. However, the foreground-background feature coupling caused by the overlapping phenomenon specific to X-ray images makes general detectors designed for natural images perform poorly. To address this issue, we propose a Category Semantic Prior Contrastive Learning (CSPCL) mechanism, which aligns the class prototypes perceived by the classifier with the content queries to correct and supplement the missing semantic information responsible for classification, thereby enhancing the model sensitivity to foreground features. To achieve this alignment, we design a specific contrastive loss, CSP loss, which includes Intra-Class Truncated Attraction (ITA) loss and Inter-Class Adaptive Repulsion (IAR) loss, and outperforms classic N-pair loss and InfoNCE loss. Specifically, ITA loss leverages class prototypes to attract intra-class category-specific content queries while preserving necessary distinctiveness. IAR loss utilizes class prototypes to adaptively repel inter-class category-specific content queries based on the similarity between class prototypes, helping disentangle features of similar categories. CSPCL is general and can be easily integrated into Deformable DETR-based models. Extensive experiments on the PIXray and OPIXray datasets demonstrate that CSPCL significantly enhances the performance of various state-of-the-art models without increasing complexity. The code will be open source once the paper is accepted.
@@ 1 Introduction
Prohibited item detection based on X-ray images is an important security inspection measure, widely used in airports, postal services, government agencies, and border control areas. With the development of computer vision technologies, researchers have started using techniques such as image classification, object detection, and semantic segmentation to assist security personnel in examining packages imaged by security screening machines for prohibited items, thereby reducing the potential risks caused by work fatigue. However, as shown in Fig. 1, X-ray images differ from natural light images in that they exhibit unique overlapping phenomena, which cause coupling of foreground and background features. This leads to the detection results of general object detection models being easily affected by background noise.
@@
There are two mainstream approaches for improving the anti-overlapping detection performance of detectors in response to the overlapping phenomenon in X-ray images. Methods based on attention mechanisms, including DAM [38], RIA [33], and FCAM [13], focus on foreground information either channel-wise or spatially. Methods based on label assignment, including IAA [16], HSS [3], and LAreg[39], provide the model with more accurate candidate boxes during the training process, allowing it to concentrate on perceiving semantic information from overlapping foreground and background features. However, the aforementioned methods are all designed for CNN-based detectors and lack a decoder structure to leverage the reliable prior knowledge provided by queries to perceive specific foreground features [2, 21, 37]. Recently, researchers have found that clarifying the classification semantic information of content queries in Deformable DETR-based models can enhance the ability of the decoder to perceive foreground features in X-ray images and improve the anti-overlapping detection ability [15, 14]. AO-DETR [15] introduces the CSA label assignment strategy, which constrains content queries to detect specific categories of objects, but its portability is relatively limited. MMCL [14] proposes a plug-and-play contrastive learning mechanism that establishes sample pairs between content queries and categorizes them according to the number of classes. This method indiscriminately repels inter-class sample pairs while attracting intra-class sample pairs, thereby clarifying the semantic information of categories. However, it lacks category prior information, which may lead to content queries not aligning with the true distribution of feature manifold of categories.
As shown in Fig. 1, the features of prohibited items such as knife, saw, and fireworks exhibit distinct differences, with varying degrees of differentiation between categories. Specifically, knife and saw have more similar color, texture, and contour information compared to knife and fireworks. Consequently, the representations of knife and saw are closer together in feature space than those of knife and fireworks. Based on this observation, we propose a novel Category Semantic Prior Contrastive Learning mechanism (CSPCL), which utilizes classifier weights as specific category prototypes to provide targeted correction and guidance to the content queries responsible for each category in the decoder layers, ensuring that the content queries align with the inherent characteristic distribution of the respective prohibited items. To achieve this multi-class multi-sample alignment, we design a Category Semantic Prior (CSP) loss, which includes Intra-Class Truncated Attraction (ITA) loss and Inter-Class Adaptive Repulsion (IAR) loss, and shows superior performance compared to classic N-pair [27] loss and InfoNCE [26] loss. Specifically, we employ ITA loss to attract content queries of the same category toward their respective category prototypes, providing the queries with category semantic prior information. The ITA loss includes a gradient truncation function that stops the attraction when the similarity between the prototypes and intra-class content queries exceeds a certain threshold, thereby preventing homogenization among content queries of the same class. Similarly, the IAR loss utilizes class prototypes to repel inter-class content queries, which has a repulsion factor adaptively adjusting the repulsion strength based on the similarity between prototype pairs, thereby facilitating sufficient differentiation between content queries of similar categories.
@@
  1. 1.
    We propose a plug-and-play CSPCL mechanism that uses contrastive learning to align category-specific content queries with the intrinsic feature distribution of the same category of prohibited items as perceived by the classifier, thereby enhancing the anti-overlapping detection capability of Deformable DETR-based models without increasing complexity.
  2. 2.
    We propose the CSP loss tailored for the multi-class multi-sample alignment task, which more effectively aligns class prototypes with content queries than classic contrastive losses like N-pair loss and InfoNCE loss.
  3. 3.
    The CSPCL mechanism demonstrates strong generalization across various Deformable DETR variants, such as RT-DETR [40], DINO [37], and AO-DETR [15], improving detection accuracy on two prohibited item datasets, including PIXray [23] and OPIXray [33].
@@ 2 Related Work
Contrastive Learning. The core idea of contrastive learning is to train more discriminative feature representations by pulling together the representations of similar samples and pushing apart the representations of dissimilar samples. Overall, contrastive learning can be divided into unsupervised contrastive learning and supervised contrastive learning. Unsupervised contrastive learning methods can be further divided into instance-wise contrastive learning and cluster-based contrastive learning. The former is represented by methods such as InstDisc [35] and SimCLR [6], where, for any given instance, its data-augmented version is treated as an intra-class sample, while other instances are considered as inter-class samples. Typically, the InfoNCE [26] loss is used for gradient updates. The latter, represented by methods such as CC [17] and TCL [18], uses pseudo-labels generated by clustering algorithms as the basis and then employs a supervised contrastive learning framework for training. Supervised contrastive learning methods have three paradigms of loss function. The first is based on the LMMN [34], which includes Triplet loss and N-pair loss. The second is based on the Softmax function, such as AM-Softmax [31], Circle loss [28], and SupCon loss [11]. The third paradigm is based on Cross-Entropy loss, including methods like EBM [7], C2AM [36], and MMCL [14]. As far as we know, the choice among these three paradigms often depends on the specific task and the reliability of the labels, with no definitive superiority among them.
@@
Prohibited Item Detector. Since the introduction of the first large-scale pseudo-color X-ray prohibited item detection dataset by SIXray [25], research on prohibited item object detection in X-ray images has become increasingly diverse and manifold. Prohibited item detectors are typically based on general object detectors and are improved to address the overlap phenomena in X-ray images. SIXray introduced the CHR mechanism to iteratively strip away overlapping background information. OPIXray [33] proposed the DOAM module to perceive the material and edge information of foreground objects. GADet [16] and Xdet [3] proposed the IAA and HSS label assignment strategies to mitigate the foreground-background class imbalance issue caused by overlapping phenomena. FDTNet [42] and FAPID [19] transform features into the frequency domain and then use self-attention mechanisms or convolutions to extract the texture and edge information of foreground objects. AO-DETR [15] proposed the first DETR-like model in the field of prohibited item detection, utilizing the CSA strategy to train category-specific content queries that are responsible for detecting specific types of contraband, thereby enhancing the ability to detect overlapping objects. However, the portability of CSA is limited. MMCL [14] proposed a plug-and-play contrastive learning strategy that helps clarify the category-specific semantic information of content queries in Deformable DETR-based models. However, it lacks prior knowledge of the feature distribution for each category, and the indiscriminate repulsion of inter-class content queries will lead to the misalignment between the feature distributions of queries and objects within the same category. This paper presents improvements to address this issue.
@@ 3.1 Explore the significance of feature alignment between objects and content queries
In this part, we examine the role of content queries in the decoder of Deformable DETR-based models and elucidate the importance of aligning object features with content queries. As shown in Fig. 2(a), using DINO as a representative model of the Deformable DETR series, given an input image, the backbone of the model first extracts multi-scale features, which are then integrated through multiple layers of encoder layers to obtain global information, commonly referred to as the encoder memory $\mathbf{M}\in\mathds{R}^{N\times C}$ [2], which is passed to the decoder. In the decoder, as depicted in Fig. 2(b), the decoder first uses the $\mathbf{M}$ to predict a set of candidate results, including classification results $\mathbf{C}\in\mathds{R}^{N\times K}$ and localization results $\mathbf{R}\in\mathds{R}^{N\times 4}$, where $K$ is the number of categories in the dataset. Then, through the query selection mechanism, the top $N_{pred}$ most reliable predictions $\mathbf{R}^{0}\in\mathds{R}^{N_{pred}\times 4}$ are filtered based on their classification confidence.
@@
$\{\mathbf{Q}_{c}^{l+1},\mathbf{R}^{l+1},\mathbf{C}^{l+1}\}=\mathcal{D}^{l}(\mathbf{Q}_{c}^{l},\mathbf{R}^{l},\mathbf{M};\theta^{l}),$ (1)
$\mathbf{R}^{l+1}=\sigma(\sigma^{-1}(\mathbf{R}^{l})+\Delta\mathbf{R}^{l+1}),$ (2)
@@
where $l\in\{x\mid x\in\mathds{Z},0\leq x<L\}$ is the decoder block index, and $L$, topically set to 6, denotes the total number of decoder blocks. The $\theta$, $\sigma(\cdot)$, $\sigma^{-1}(\cdot)$, and $\mathcal{H}_{R,C}$ denote learnable parameters, sigmoid, inverse sigmoid [41], and prediction heads of regression and classification [2], respectively. Equation 3 shows that content queries almost directly determine the prediction results at each layer of the decoder. To analyze the update process of content queries more deeply, we further decompose the decoder in Eq. 1 into self-attention mechanisms and deformable attention mechanisms for separate explanations.
@@
$\mathbf{Q}_{s,c}^{l}=\mathcal{D}_{s}^{l}(\mathbf{Q}_{s}^{l},\mathbf{K}_{s}^{l},\mathbf{V}_{s}^{l};\theta_{s}^{l}),$ (4)
$\mathbf{Q}_{s}^{l}=\mathbf{K}_{s}^{l}=\mathbf{Q}_{c}^{l}+\mathbf{Q}_{p}^{l},\quad\mathbf{V}_{s}^{l}=\mathbf{Q}_{c}^{l},$ (5)
@@
$\mathbf{Q}_{c}^{l+1}=\mathbf{Q}_{d,c}^{l+1}=\mathcal{D}_{d}^{l}(\mathbf{Q}_{d}^{l},\mathbf{R}^{l},\mathbf{M};\theta_{d}^{l})$ (6)
$\mathbf{Q}_{d}^{l}=\mathbf{Q}_{s,c}^{l}+\mathbf{Q}_{p}^{l}.$ (7)
@@ 3.2 Category semantic prior contrastive learning (CSPCL)
We propose the CSPCL mechanism, as illustrated in Fig. 2(c), which supplements and corrects the category semantic information of content queries by aligning them with the class prototypes perceived by the classifier weights. In the preprocessing stage, we group the content queries $\mathbf{Q}_{c}^{l}\in\mathds{R}^{N_{pred}\times M}$ from the $l$-th layer based on the number of categories $K$, resulting in $K$ groups of category-specific content queries $\mathbf{Q}_{c,k}^{l}\in\mathds{R}^{n\times M}$, where $k\in\{x\mid x\in\mathds{Z},0<x\leq K\}$, and $n$ is the quotient of the number of content queries $N_{pred}$ divided by the number of categories $K$. Next, the classifier weights $\mathbf{W}\in\mathds{R}^{K\times M}$ are dimensionally expanded and adjusted to align with the dimensions of $\mathbf{Q}_{c}^{l}$, resulting in prototypes $\mathbf{P}\in\mathds{R}^{N_{pred}\times M}$. In this way, the prototype, $\mathbf{P}_{k}\in\mathds{R}^{n\times M}$, which is responsible for category $k$, shares the same dimensions as $\mathbf{Q}_{c,k}^{l}$. The prototypes $\mathbf{P}_{k}$ and the content queries $\mathbf{Q}_{c,k}^{l}$ consist of $n$ sample pairs for each category $k$. For this multi-class multi-sample alignment task, we specifically designed the CSP loss consisting of the Intra-Class Truncated Attraction (ITA) loss and the Inter-Class Adaptive Repulsion (IAR) loss. The ITA loss attracts the prototypes to their corresponding intra-class content queries, while the IAR loss repels the prototypes from inter-class content queries. The formula for CSP loss is as follows:
@@
ITA loss. To attract category-specific content queries to the feature manifold of intra-class prohibited items, a simple and effective method is to compute the cross-entropy loss using the cosine similarity between category prototypes and intra-class category-specific content queries. However, even prohibited items of the same category still exhibit feature variations under different backgrounds and poses, meaning that intra-class category-specific content queries need to maintain a certain degree of variability to align with these differences, thereby improving the detection accuracy. Therefore, we specifically designed a gradient truncation mechanism to ensure that content queries retain enough variance. The overall formula for the ITA loss is as follows:
$\begin{split}\mathcal{L}_{ITA}(\mathbf{P,Q},K)=&\frac{-1}{Kn(n-1)}\sum_{k=1}^{K}\sum_{i=1}^{n}\sum_{j=1}^{n}\\ &\text{log}\Big{(}\mathcal{T}\big{(}\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k}),\gamma\big{)}\Big{)},\end{split}$ (9)
$\mathcal{T}(x,\gamma)=\begin{cases}1-\gamma,&1-\gamma<x\leq 1\\ x,&0<x<1-\gamma\\ 0,&-1\leq x<0\end{cases}$ (10)
@@
where $\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k})$ is the cosine similarity between the $i$-th prototype of the $k$-th class and $j$-th content query of the $k$-th class. $\mathcal{T}(x,\gamma)$ is the gradient truncation function. As shown in Fig. 3, when the input value $x$, i.e. $\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k})$, exceeds the threshold $1-\gamma$, the returned ITA loss value is constant $1-\gamma$, and its derivative becomes 00. At this point, the prototype stops attracting the intra-class content queries, allowing the content queries to retain the necessary variability and margin among queries. Therefore, by adjusting the value of $\gamma$, we can control the variance of similarities of intra-class content queries.
Figure 3: (a) and (b) are the curves of the ITA loss and its derivative, respectively.
@@
Figure 4: The 3D surface plots of our IAR loss with different $\tau$. The samples A and B have the same $\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{q}_{j}^{k_{2}})$, but $\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{p}_{j}^{k_{2}})$ of A is smaller that of B. When $\tau=0.1$, the IAR loss values of A and B are almost the same, but when $\tau=1$, The loss value of A is much smaller than that of B.
@@
IAR loss. Although the ITA attracts content queries toward the intra-class prototypes, some stubborn samples still drift away from the sample center, and may even appear in the feature manifold of other categories. A simple solution is to use the prototype to repel inter-class content queries indiscriminately. However, the differences between the prototypes of different categories vary naturally. As shown in Fig. 1, the feature manifolds of knife and saw have less margin than those of knife and firework, which means the former is naturally more coupled than the latter, requiring stronger repulsion by the prototype to create enough margin for clarifying the semantic information of content queries for better detection performance [1, 10, 15, 14]. Therefore, we propose the IAR loss as follows:
$\begin{split}\mathcal{L}_{IAR}(\mathbf{P,Q},K)=&\frac{-1}{K(K-1)n^{2}}\sum_{k_{1}=1}^{K}\sum_{k_{2}=1}^{K}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathds{1}\left[k_{1}\neq k_{2}\right]\\ &\mathcal{R}(k_{1},k_{2})\text{log}(1-\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{q}_{j}^{k_{2}})),\end{split}$ (11)

### v2

@@ Abstract
Prohibited item detection based on X-ray images is one of the most effective security inspection methods. However, the foreground-background feature coupling caused by the overlapping phenomenon specific to X-ray images makes general detectors designed for natural images perform poorly. To address this issue, we propose a Category Semantic Prior Contrastive Learning (CSPCL) mechanism, which aligns the class prototypes perceived by the classifier with the content queries to correct and supplement the missing semantic information responsible for classification, thereby enhancing the model sensitivity to foreground features. To achieve this alignment, we design a specific contrastive loss, CSP loss, which comprises the Intra-Class Truncated Attraction (ITA) loss and the Inter-Class Adaptive Repulsion (IAR) loss, and outperforms classic contrastive losses. Specifically, the ITA loss leverages class prototypes to attract intra-class content queries and preserves essential intra-class diversity via a gradient truncation function. The IAR loss employs class prototypes to adaptively repel inter-class content queries, with the repulsion strength scaled by prototype-prototype similarity, thereby improving inter-class discriminability, especially among similar categories. CSPCL is general and can be easily integrated into Deformable DETR-based models. Extensive experiments on the PIXray, OPIXray, PIDray, and CLCXray datasets demonstrate that CSPCL significantly enhances the performance of various state-of-the-art models without increasing inference complexity. The code is publicly available at https://github.com/Limingyuan001/CSPCL.
@@ 1 Introduction
Prohibited item detection based on X-ray images is an important security inspection measure, widely used in airports, postal services, government agencies, and border control areas. With the development of computer vision technologies, researchers have started using techniques such as image classification, object detection, and semantic segmentation to assist security personnel in examining packages imaged by security screening machines for prohibited items, thereby reducing the potential risks caused by work fatigue. However, as shown in Figure˜1, X-ray images differ from natural light images in that they exhibit unique overlapping phenomena, which cause coupling of foreground and background features. This leads to the detection results of general object detection models being easily affected by background noise.
There are two mainstream approaches for improving the anti-overlapping detection performance of detectors in response to the overlapping phenomenon in X-ray images. Methods based on attention mechanisms, including DAM [1], RIA [2], and FCAM [3], focus on foreground information either channel-wise or spatially. Methods based on label assignment, including IAA [4], HSS [5], and LAreg[6], provide the model with more accurate candidate boxes during the training process, allowing it to concentrate on perceiving semantic information from overlapping foreground and background features. However, the aforementioned methods are all designed for CNN-based detectors and lack a decoder structure to leverage the reliable prior knowledge provided by queries to perceive specific foreground features [7; 8; 9]. Recently, researchers have found that clarifying the classification semantic information of content queries in Deformable DETR-based models can enhance the ability of the decoder to perceive foreground features in X-ray images and improve the anti-overlapping detection ability [10; 11]. AO-DETR [10] introduces the CSA label assignment strategy, which constrains content queries to detect specific categories of objects, but its portability is relatively limited. MMCL [11] proposes a plug-and-play contrastive learning mechanism that establishes sample pairs between content queries and categorizes them according to the number of classes. This method indiscriminately repels inter-class sample pairs while attracting intra-class sample pairs, thereby clarifying the semantic information of categories. However, it lacks category prior information, which may lead to content queries not aligning with the true distribution of feature manifold of categories.
@@
As shown in Figure˜1, the features of prohibited items such as knife, saw, and fireworks exhibit distinct differences, with varying degrees of differentiation between categories. Specifically, knife and saw have more similar color, texture, and contour information compared to knife and fireworks. Consequently, the representations of knife and saw are closer together in feature space than those of knife and fireworks. Based on this observation, we propose a novel Category Semantic Prior Contrastive Learning mechanism (CSPCL), which utilizes classifier weights as specific category prototypes to provide targeted correction and guidance to the content queries responsible for each category in the decoder layers, ensuring that the content queries align with the inherent characteristic distribution of the respective prohibited items. To achieve this multi-class multi-sample alignment, we design a contrastive loss, called the Category Semantic Prior (CSP) loss, which comprises the Intra-Class Truncated Attraction (ITA) loss and the Inter-Class Adaptive Repulsion (IAR) loss, and shows superior performance compared to classic contrastive losses [12; 13]. Specifically, we employ the ITA loss to attract content queries of the same category toward their respective category prototypes, providing the queries with category semantic prior information. It provides a gradient truncation function that stops the attraction when the similarity between the prototypes and intra-class content queries exceeds a certain threshold, thereby preventing the homogenization of intra-class content queries. Inversely, the IAR loss utilizes class prototypes to repel inter-class content queries, which has a repulsion factor adaptively adjusting the repulsion strength based on the similarity between prototype pairs, thereby facilitating inter-class content queries to learn discriminative identifying features of similar categories, and obtain adequate separability.
@@
  • We propose a plug-and-play CSPCL mechanism that uses contrastive learning to align content queries with the intrinsic feature distribution of the same category of prohibited items as perceived by the classifier, thereby enhancing the anti-overlapping detection capability of Deformable DETR-based models without increasing inference complexity.
  • We propose the CSP loss tailored for the multi-class multi-sample alignment task, which more effectively aligns class prototypes with content queries than classic contrastive losses.
  • The CSPCL mechanism demonstrates strong generalization across various Deformable DETR variants, such as RT-DETR [14], DINO [9], and AO-DETR [10], improving detection accuracy on two prohibited item datasets, including PIXray [15] and OPIXray [2]. Furthermore, we validate CSPCL on the large-scale prohibited item detection datasets PIDray [1] and CLCXray [6], further advancing the state-of-the-art standards.
@@ 2 Related Work
Contrastive Learning. The core idea of contrastive learning is to train more discriminative feature representations by pulling together the representations of similar samples and pushing apart the representations of dissimilar samples. Overall, contrastive learning can be divided into unsupervised contrastive learning and supervised contrastive learning. Unsupervised contrastive learning methods can be further divided into instance-wise contrastive learning and cluster-based contrastive learning. The former is represented by methods such as InstDisc [24] and SimCLR [25], where, for any given instance, its data-augmented version is treated as an intra-class sample, while other instances are considered as inter-class samples. Typically, the InfoNCE [13] loss is used for gradient updates. The latter, represented by methods such as CC [26] and TCL [27], uses pseudo-labels generated by clustering algorithms as the basis and then employs a supervised contrastive learning framework for training. Supervised contrastive learning methods have three paradigms of loss function. The first is based on the LMMN [28], which includes Triplet loss and N-pair loss. The second is based on the Softmax function, such as AM-Softmax [29], Circle loss [30], and SupCon loss [31]. The third paradigm is based on Cross-Entropy loss, including methods like EBM [32], C2AM [33], and MMCL [11]. As far as we know, the choice among these three paradigms often depends on the specific task and the reliability of the labels, with no definitive superiority among them. In this paper, we propose a targeted contrastive loss function, CSP loss, for the task of guiding content queries with category semantic priors.
@@
Prohibited Item Detector. Since the introduction of the first large-scale pseudo-color X-ray prohibited item detection dataset by SIXray [34], research on prohibited item object detection in X-ray images has become increasingly diverse and manifold. Prohibited item detectors are typically based on general object detectors and are improved to address the overlap phenomena in X-ray images [35; 36; 37; 4; 38; 39; 40]. SIXray introduced the CHR mechanism to iteratively strip away overlapping background information. OPIXray [2] proposed the DOAM module to perceive the material and edge information of foreground objects. GADet [4] and Xdet [5] proposed the IAA and HSS label assignment strategies to mitigate the foreground-background class imbalance issue caused by overlapping phenomena. FDTNet [41] and FAPID [42] transform features into the frequency domain and then use self-attention mechanisms or convolutions to extract the texture and edge information of foreground objects. AO-DETR [10] proposed the first DETR-like model in the field of prohibited item detection, utilizing the CSA strategy to train category-specific content queries that are responsible for detecting specific types of contraband, thereby enhancing the ability to detect overlapping objects. However, the portability of CSA is limited. MMCL [11] proposed a plug-and-play contrastive learning strategy that helps clarify the category-specific semantic information of content queries in Deformable DETR-based models. However, it lacks prior knowledge of the feature distribution for each category, and the indiscriminate repulsion of inter-class content queries will lead to the misalignment between the feature distributions of queries and objects within the same category. This paper presents improvements to address this issue.
@@ 3.1 Explore the significance of feature alignment between objects and content queries
In this part, we examine the role of content queries in the decoder of Deformable DETR-based models and elucidate the importance of aligning object features with content queries. As shown in Figure˜2(a), using DINO as a representative model of the Deformable DETR series, given an input image, the backbone of the model first extracts multi-scale features, which are then integrated through multiple layers of encoder layers to obtain global information, commonly referred to as the encoder memory $\mathbf{M}\in\mathds{R}^{N\times C}$ [7], which is passed to the decoder. In the decoder, as depicted in Figure˜2(b), the decoder first uses the $\mathbf{M}$ to predict a set of candidate results, including classification results $\mathbf{C}\in\mathds{R}^{N\times K}$ and localization results $\mathbf{R}\in\mathds{R}^{N\times 4}$, where $K$ is the number of categories in the dataset. Then, through the query selection mechanism, the top $N_{pred}$ most reliable predictions $\mathbf{R}^{0}\in\mathds{R}^{N_{pred}\times 4}$ are filtered based on their classification confidence.
@@
$\{\mathbf{Q}_{c}^{l+1},\mathbf{R}^{l+1},\mathbf{C}^{l+1}\}=\mathcal{D}^{l}(\mathbf{Q}_{c}^{l},\mathbf{R}^{l},\mathbf{M};\theta^{l}),\quad\mathbf{R}^{l+1}=\sigma(\sigma^{-1}(\mathbf{R}^{l})+\Delta\mathbf{R}^{l+1}),$ (1)
@@
where $l\in\{x\mid x\in\mathds{Z},0\leq x<L\}$ is the decoder block index, and $L$, topically set to 6, denotes the total number of decoder blocks. The $\theta$, $\sigma(\cdot)$, $\sigma^{-1}(\cdot)$, and $\mathcal{H}_{R,C}$ denote learnable parameters, sigmoid, inverse sigmoid [16], and prediction heads of regression and classification [7], respectively. Eq.˜2 shows that content queries almost directly determine the prediction results at each layer of the decoder. To analyze the update process of content queries more deeply, we further decompose the decoder in Eq.˜1 into self-attention mechanisms and deformable attention mechanisms for separate explanations.
@@
$\mathbf{Q}_{s,c}^{l}=\mathcal{D}_{s}^{l}(\mathbf{Q}_{s}^{l},\mathbf{K}_{s}^{l},\mathbf{V}_{s}^{l};\theta_{s}^{l}),\quad\mathbf{Q}_{s}^{l}=\mathbf{K}_{s}^{l}=\mathbf{Q}_{c}^{l}+\mathbf{Q}_{p}^{l},\quad\mathbf{V}_{s}^{l}=\mathbf{Q}_{c}^{l},$ (3)
@@
$\mathbf{Q}_{c}^{l+1}=\mathbf{Q}_{d,c}^{l+1}=\mathcal{D}_{d}^{l}(\mathbf{Q}_{d}^{l},\mathbf{R}^{l},\mathbf{M};\theta_{d}^{l}),\quad\mathbf{Q}_{d}^{l}=\mathbf{Q}_{s,c}^{l}+\mathbf{Q}_{p}^{l}.$ (4)
@@ 3.2 Category semantic prior contrastive learning (CSPCL)
We propose the CSPCL mechanism, as illustrated in Figure˜2(c), which supplements and corrects the category semantic information of content queries by aligning them with the class prototypes perceived by the classifier weights. In the preprocessing stage, we group the content queries $\mathbf{Q}_{c}^{l}\in\mathds{R}^{N_{pred}\times M}$ from the $l$-th layer based on the number of categories $K$, resulting in $K$ groups of category-specific content queries $\mathbf{Q}_{c,k}^{l}\in\mathds{R}^{n\times M}$, where $k\in\{x\mid x\in\mathds{Z},0<x\leq K\}$, and $n$ is the quotient of the number of content queries $N_{pred}$ divided by the number of categories $K$. Next, the classifier weights $\mathbf{W}\in\mathds{R}^{K\times M}$ are dimensionally expanded and adjusted to align with the dimensions of $\mathbf{Q}_{c}^{l}$, resulting in prototypes $\mathbf{P}\in\mathds{R}^{N_{pred}\times M}$. In this way, the prototype, $\mathbf{P}_{k}\in\mathds{R}^{n\times M}$, which is responsible for category $k$, shares the same dimensions as $\mathbf{Q}_{c,k}^{l}$. The prototypes $\mathbf{P}_{k}$ and the content queries $\mathbf{Q}_{c,k}^{l}$ consist of $n$ sample pairs for each category $k$. For this multi-class multi-sample alignment task, we specifically designed the CSP loss consisting of the Intra-Class Truncated Attraction (ITA) loss and the Inter-Class Adaptive Repulsion (IAR) loss. The IAR loss employs category prototypes to repel inter-class content queries, while the ITA loss attracts content queries towards the corresponding intra-class prototypes, thereby enhancing inter-class discriminability—especially among similar categories—while preserving essential intra-class diversity. The formula for CSP loss is as follows:
@@
ITA loss. To attract category-specific content queries to the feature manifold of intra-class prohibited items, a simple and effective method is to compute the cross-entropy loss using the cosine similarity between category prototypes and intra-class content queries. However, even prohibited items of the same category still exhibit feature variations under different backgrounds and poses, meaning that intra-class content queries need to maintain a certain degree of variability to align with these differences, thereby improving the detection accuracy. Therefore, we specifically designed a gradient truncation mechanism to ensure that content queries retain enough variance. The overall formula for the ITA loss is as follows:
$\mathcal{L}_{ITA}(\mathbf{P,Q},K)=\frac{-1}{Kn(n-1)}\sum_{k=1}^{K}\sum_{i=1}^{n}\sum_{j=1}^{n}\text{log}\Big(\mathcal{T}\big(\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k}),\gamma\big)\Big),\mathcal{T}(x,\gamma)=\begin{cases}1-\gamma,&1-\gamma<x\leq 1\\ x,&0<x<1-\gamma\\ 0,&-1\leq x<0\end{cases}$ (6)
@@
where $\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k})$ is the cosine similarity between the $i$-th prototype of the $k$-th class and $j$-th content query of the $k$-th class. $\mathcal{T}(x,\gamma)$ is the gradient truncation function. As shown in Figure˜3, when the input value $x$, i.e. $\text{sim}(\mathbf{p}_{i}^{k},\mathbf{q}_{j}^{k})$, exceeds the threshold $1-\gamma$, the returned ITA loss value is constant $1-\gamma$, and its derivative becomes 0. At this point, the prototype stops attracting the intra-class content queries, allowing the content queries to retain the necessary variability and margin among queries. Therefore, by adjusting the value of $\gamma$, we can control the variance (diversity) of similarities of intra-class content queries.
@@
Figure 3: (a) and (b) are the curves of the ITA loss and its derivative, respectively. (c) and (d) are the 3D surface plots of our IAR loss with different $\tau$. The samples A and B have the same $\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{q}_{j}^{k_{2}})$, but $\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{p}_{j}^{k_{2}})$ of A is smaller that of B. When $\tau=0.1$, the IAR loss values of A and B are almost the same, but when $\tau=1$, The loss value of A is much smaller than that of B.
@@
IAR loss. Although the ITA attracts content queries toward the intra-class prototypes, some stubborn samples still drift away from the sample center, and may even appear in the feature manifold of other categories. A simple solution is to use the prototype to repel inter-class content queries indiscriminately. However, the differences between the prototypes of different categories vary naturally. As shown in Figure˜1, the feature manifolds of knife and saw have less margin than those of knife and firework, which means the former is naturally more coupled than the latter, requiring stronger repulsion by the prototype to create enough margin for clarifying the semantic information of content queries for better detection performance [44; 45; 10; 11]. Therefore, we propose the IAR loss as follows:
$\mathcal{L}_{IAR}(\mathbf{P,Q},K)=\frac{-1}{K(K-1)n^{2}}\sum_{k_{1}=1}^{K}\sum_{k_{2}=1}^{K}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathds{1}\left[k_{1}\neq k_{2}\right]\\ \mathcal{R}(k_{1},k_{2})\text{log}(1-\text{sim}(\mathbf{p}_{i}^{k_{1}},\mathbf{q}_{j}^{k_{2}})),$ (7)